### v1

@@ 1 Introduction
An alternative approach to compact on-chip optical devices with good in-situ tunability over the emission characteristics is realised with optical lattices of non-equilibrium exciton-polariton Bose-Einstein condensates [17, 18] which, due to their strong nonlinearities and optically tunable ballistic coupling mechanism [19], are being explored as promising reconfigurable elements for all-optical computing [20, 21], spin simulation [4, 22], angular momentum generation [23, 24], topologically protected lasing [25, 26, 27], and more. Although many of the aforementioned works have been conducted at cryogenic temperatures in II-VI or III-V semiconductor quantum well microcavities, increasing effort has been shifted towards more application-friendly room temperature materials [28, 29, 30], in particular organic materials relevant to the current study [31, 32, 33]. However, for room temperature organic-based light-emission devices, the coherent emission usually appears perpendicular to the cavity surface (at $k_{||}=0$) and therefore coupling between the emission nodes (i.e., pumped condensates) is hard to achieve and usually requires significant overlap between the condensate wave-functions [31, 32, 34]. Besides, the requirement for strong light-matter coupling imposes strict conditions on the parameters of both photonic and excitonic components of the system, restricting tuneability and scalability. In this work we attempt to fill the gap between the optically reconfigurable ballistic lattices of polaritonic condensates and tuneable organic microcavities operating at room temperature in the weak light-matter coupling regime, where in-plane phase-locking between separated optically excited lasing states has yet not been reported.
@@ 2.1 Platform design and approach
Figure 1: Electrically controlled dyad lasing supermode and the blueshift-induced in-plane momentum of coherent emission from a single lasing state. (a,b) Schematic showing the microcavity filled with liquid crystals and P580 laser dye pumped in a dyad configuration with (a) phase-locked spatially extended supermode without external voltage applied and (b) two uncoupled lasing states at 1600 mV voltage correspondingly. Emission characteristics of a single lasing state: (c-e) (above) energy-resolved momentum-space along $k_{y}=0$ for three values of external voltage, corresponding to (c) 0 mV, (d) 1600 mV, and (e) 1840 mV, pumped below ($0.5P_{th}$, left) and above ($1.5P_{th}$, right) the lasing threshold for each case and (below) momentum-space pumped above ($1.5P_{th}$) the lasing threshold, dot-dashed and dotted curves show (c-d) horisontally and vertically polarised or (e) right-circularly and left-circularly polarised cavity modes; (f) momentum-space distribution of the normalised $S_{1}$ (top and middle) and $S_{3}$ (bottom) Stokes components corresponding to momentum-space at (c-e); (g) photoluminescence intensity (black circles), full-width at half-maximum (FWHM, red squares), and (h) energy of the lasing state versus pump power at zero voltage, extracted from the fitting of the emission spectral profile obtained by integrating dispersion data around $k_{x}=0$ over ±$\pm 0.2\mu m^{-1}$ of the energy-resolved momentum space measurements. For better visibility energy resolved images (c-e) are illustrated in logarithmic scale saturated below 0.05.
@@
Figure 1(g,h) presents the dependence of PL intensity, FWHM, and emission energy on the pumping power at 0 mV external voltage. Initially, a linear increase in intensity is observed, followed by a sharp nonlinear behaviour accompanied by a step-like drop in FWHM, both hallmarks of a transition from spontaneous emission to photonic lasing. This transition is further evidenced by a sharp blueshift of the emission relative to the cavity dispersion minimum below threshold. In conventional inorganic systems, such as GaAs-quantum well microcavities with Wannier–Mott excitons, such blueshift near threshold is a recognized signature of polariton lasing [42]. However, this interpretation does not straightforwardly apply to organic microcavities with Frenkel excitons, which are highly localized and do not support comparable repulsive interactions. It has been shown [43, 44] that, in organic systems, the blueshift primarily arises from the saturation of molecular transitions and the associated refractive index change and hence, is present in both weak and strong light-matter coupling regimes and, importantly, is wavelength dependent, see Fig 1(c,d,e). Under tight optical pumping, it creates a localized potential sufficient to induce the outflow of coherent photons with finite average absolute value of in-plane momentum ($k_{||}\neq 0$), resulting in ring-shaped emission in momentum-space (or two split rings relative to $k_{x}=0$ in the RD regime), with the radius $k_{l}$ determined by the amount of pump-induced blueshift [see momentum-space PL at the bottom of Figs. 1(c,d,e)]. This ring is similar to resonant Rayleigh scattering ring when waves are scattering of disorder and couple to many different momenta [45]. Moreover, for tuneable LCMCs, this blueshift depends on the spectral detuning between the P580 dye absorption and the cavity mode, which can be modulated via applied voltage [Fig. 1(c,d)]. Consequently, the $k$-space profile of the coherent emission is electrically controllable through the modulation of the local potential. While the underlying mechanisms differ between room-temperature organic microcavities and cryogenically operated GaAs-based systems, both generate coherent emission with finite in-plane momentum, which is a key requirement for mutual phase-locking between lasing sites and the emergence of spatially extended lasing "supermodes" via in-plane coherent energy exchange.
@@ 2.2 Dyad lasing supermode
Figure 2: Dyad supermode lasing state. (a-i) Experimental images of the dyad supermode lasing state emission in (a-c) real-space, (d-f) momentum space, and energy-resolved momentum-space along $k_{y}=0$ (g,h,i) for two pump spots (black circles) separated by (a,d,g) 9 $\mu$m, (b,e,h) 11 $\mu$m, and (c,f,i) 13 $\mu$m. (j,k) Numerically simulated (j) real-space and (k) momentum-space distribution of the supermode lasing state emission corresponding to the experimental data at (b,e). (l) Dependence of the extended lasing state mode number n on the pump separation distance. Each spot was pumped with the fixed spot size Gaussian beam at $P_{1,2}=1.5P^{(1)}_{th}$, where $P^{(1)}_{th}$ is the lasing threshold for a single isolated lasing state. For better visibility real-space images (a-c,j) are illustrated in logarithmic scale saturated below 0.01.
@@
Fig. 2(a-f) shows the experimentally measured real- (a-c) and momentum- (d-f) space distributions of the PL from the spatially extended supermode lasing state formed due to the mutual exchange of coherent photons out-flowing from each of the two individually pumped lasing spots. In the absence of coherent energy exchange one would observe two independent lasing spots in the real space with the PL intensity monotonically decaying outside of the pumping spot areas, Fig. 1(b), while Fourier-space would have a ring-shaped pattern similar to the single state case. Instead, here we observe phase-locking between two individually pumped lasing states, evidenced by a characteristic oscillatory interference pattern of PL intensity in both real and Fourier spaces along the axis connecting the pump spots. This interference pattern directly reveals a macroscopic coherent wavefunction extending over distances up to $d=20\mu\mathrm{m}$. The resulting dyad lasing supermodes exhibit either even or odd parity with respect to their reflection axis at $x=0$. The parity is determined by the relative phase between the lasing centres ($\Delta\phi=0$ or $\pi$), depending on the separation $d$ between the pump spots, as expected and observed previously for polariton dyads [46]. Maximum intensity at the midpoint indicates in-phase synchronisation ($\Delta\phi=0$), while a minimum of the intensity in the centre reflects anti-phase synchronisation ($\Delta\phi=\pi$). In both cases, a single mode of corresponding parity was present in energy-resolved Fourier-space PL [Fig. 2(g-i)]. We quantitatively characterise the supermode by introducing the mode number $n$ corresponding to the number of nodes between the pump spots analogous to the mode numbers used to describe the higher-order Hermite-Gaussian transverse modes. A linear increase in the mode number $n$ of the spatially extended supermode is visible when the distance between pumping spots $d$ is increased, Fig. 2(l).
@@
To theoretically reproduce the real and Fourier-space PL from the dyad lasing state, we perform numerical simulations based on a two-dimensional (2D) non-Hermitian Schrödinger equation under drive and dissipation (a simplified version of Maxwell-Bloch equations [47, 48]). Details about the theoretical model employed are presented in Methods (Sec. 4.3). In Fig. 2(j,k) we show the simulation results for the dyad lasing state separated by $d=11\>\mu\mathrm{m}$ corresponding to the experimental data presented in Fig. 2(b,e). Corresponding simulations of the lasing states for $d=9\>\mu$m and $13\>\mu$m can be found in the SI. The slight mismatch between the experimental and numerically obtained data is due to the local disorder morphology of the sample and is mostly present in real-space PL profiles.
@@
Amazingly, in Fig. 4(c,g) we observe revival of the dyad coupling when entering the RD SOC regime at higher voltages $V=1840$ mV. Here, mixing between two distinct cavity modes results in a dispersion relation made of two off-centred cross-circularly polarized parabolas or valleys, displaced along $k_{x}$ axis parallel to the axis of the dyad [see Eq. (1) and Fig.1(e)]. Since the bottom of each dispersion branch remains similar in energy to that of the vertically polarised cavity mode at $V=0$ mV, the wavelength-dependent PL blueshift above the threshold is unchanged compared to the initial case [Fig.1(c,e)], as is the corresponding ring radius in each valley in Fourier space [Fig.4(e)]. Here, we observe emission from two separate ballistic rings in Fourier space displaced symmetrically along $k_{x}$ with opposite circular polarisation corresponding to iso-energy curves of each paraboloid (see the inset in Fig. 4(g) for $S_{3}(\mathbf{k}_{\parallel})$ Stokes parameter distribution in momentum space). We note that due to RD SOC this regime demonstrates spin-dependent anisotropy of the supermode emission with $\langle k_{x}^{\sigma+}\rangle>0>\langle k_{x}^{\sigma-}\rangle$ and $\langle k_{y}^{\sigma+}\rangle=\langle k_{y}^{\sigma-}\rangle=0$ with a clear fringe pattern observed in Fourier space, proving the phase-locking revival between the two pumped lasing states. Phase-locking revival was also observed in restoring of spatial coherence in the interferometry measurements, see in SI. In real space the intensity of the fringes between the pump spots is lowered relative to the case of $V=0$ mV. We attribute this to the local inhomogeneity affecting the real space distribution due to the reorganisation of LCs under external voltage.
@@
where $m_{x,y}$ are renormalized effective photon masses along the $x$ and $y$ directions [35] and $\delta$ describes splitting between horizontal and vertical linear polarized modes at $k_{\parallel}=0$. The modified dispersion leads to a higher photon population in the mode with linear polarisation component parallel to the pump polarization. The emission still partially occupies two ballistic circles in the momentum space PL as seen at high wavenumbers in Fig. 4(h). The radius of the circles is larger since the emission energy is now located further above the minimum of the dispersion (2). Interestingly, here the dyad mostly populates the low momentum intersect between the two circles in momentum space corresponding to photons with mostly linear vertical polarization projection. The result is a peculiar distribution of the emission in Fourier space in which the emission anisotropy is reversed relative to the case of pure RD SOC regime ($\langle|k_{y}|\rangle>\langle|k_{x}|\rangle$). Control of the in-plane momentum anisotropy is another important feature of our system as it potentially offers a way for the realisation of a direction-dependent coupling in two-dimensional arrays of phase-locked coherent emitters. Similar to the previous case the fringe pattern in real space is distorted and observable only around pumping spots. However, it is clearly visible in the Fourier space over the whole area of emission distribution. We verify our interpretation through 2D non-Hermitian Schrödinger (paraxial Maxwell-Bloch) simulations shown in Fig. 5, see Sec. 4.3 for details.
@@ 2.4 Unconventional coupling in 1D lasing supermode
In coherently coupled photonic systems the nearest-neighbour (NN) coupling dominates over the next-nearest-neighbour (NNN), including systems based on coupled lasers [49], VCSELs arrays, polaritonic and photonic lattices [3]. This is easily understandable, as the coupling strength is usually defined by the mutual overlap of the coupled states wavefunctions, which decay exponentially with the distance from each state centre. Overcoming this behaviour is of prior importance for universal applications of coherently coupled photonic arrays as all-optical simulators  [50]. The possibility of NNN coupling with polaritonic condensates in GaAs-based microcavities at cryogenic temperatures has been shown previously using distance-dependant spin-screening of NN interaction leveraging the optical spin Hall effect [51]. Same approach based on the TE-TM field is not straightforwardly applicable in organic microcavities, but it can be adapted, because the polarisation basis in these systems is defined by horizontal and vertical components, following the orientation basis of the molecular dipole moment. As we show above, the polarisation of the lasing state is then inherited from the pump polarisation and is linear as long as SOC remains small, see Sec. 2.1. As a result, the efficient coupling between the lasing spots is only possible while they share the same linear polarisation inherited from the pumps. In Fig. 6 we demonstrate the realisation of NNN coupling regime in a chain of 3 coupled lasing states, by measuring the real space PL (a-c), momentum space PL (d-f), and (g-i) real space $S_{1}$ Stokes parameter of the non-normalized Stokes vector of the extended lasing state emission [see Eq. (3)]. The corresponding numerical simulations obtained via the 2D non-Hermitian Schrödinger model are shown in the Sec. 4.3.
@@
However, when we flip the polarisation of the central pump spot to the vertical one, both real and Fourier-space distributions change dramatically, see Fig. 6(c,f,i). First, the fringes in the real space are not visible, due to the low intensity similar to the case of the far-separated dyad [Fig. 6(a)]. The Fourier space shows a superposition of two states. The first originates from a NNN type of coupling between the edge states which is visible with the odd parity and the same periodicity as the one observed for the far-separated dyad. The second mode is ring-shaped and corresponds to the Fourier space profile of an isolated central lasing state. Thus we eliminate the coupling between the nearest neighbours by exciting the lasing state in the middle with the opposite polarisation to the outer pump spots, while the NNN coupling is still present.

### v2

@@ 1 Introduction
An alternative approach to compact on-chip optical devices with good in-situ tunability over the emission characteristics is realised with optical lattices of non-equilibrium exciton-polariton Bose-Einstein condensates [17, 18] which, due to their strong nonlinearities and optically tunable ballistic coupling mechanism [19], are being explored as promising reconfigurable elements for all-optical computing [20, 21], spin simulation [4, 22], angular momentum generation [23, 24], topologically protected lasing [25, 26, 27], and more. Although many of the aforementioned works have been conducted at cryogenic temperatures in II-VI or III-V semiconductor quantum well microcavities, increasing effort has been shifted towards more application-friendly room temperature materials [28, 29, 30], in particular organic materials relevant to the current study [31, 32, 33]. However, for room temperature organic-based light-emission devices, the coherent emission usually appears perpendicular to the cavity surface (at $k_{||}=0$) and therefore coupling between the emission nodes (i.e., pumped condensates) is hard to achieve and usually requires significant overlap between the condensate wave-functions [31, 32, 34, 35, 36]. Besides, the requirement for strong light-matter coupling imposes strict conditions on the parameters of both photonic and excitonic components of the system, restricting tuneability and scalability. In this work we attempt to fill the gap between the optically reconfigurable ballistic lattices of polaritonic condensates and tuneable organic microcavities operating at room temperature in the weak light-matter coupling regime, where in-plane phase-locking between separated optically excited lasing states has yet not been reported.
@@ 2.1 Platform design and approach
Figure 1: Electrically controlled dyad lasing supermode and the blueshift-induced in-plane momentum of coherent emission from a single lasing state. (a,b) Schematic showing the microcavity filled with liquid crystals and P580 laser dye pumped in a dyad configuration with (a) phase-locked spatially extended supermode without external voltage applied and (b) two uncoupled lasing states at 1600 mV voltage correspondingly. Emission characteristics of a single lasing state: (c-e) (above) energy-resolved momentum-space along $k_{y}=0$ for three values of external voltage, corresponding to (c) 0 mV, (d) 1600 mV, and (e) 1840 mV, pumped below ($0.5P_{th}$, left) and above ($1.5P_{th}$, right) the lasing threshold for each case and (below) momentum-space pumped above ($1.5P_{th}$) the lasing threshold, dot-dashed and dotted curves show (c-d) horisontally and vertically polarised or (e) right-circularly and left-circularly polarised cavity modes; (f) momentum-space distribution of the normalised $S_{1}$ (top and middle) and $S_{3}$ (bottom) Stokes components corresponding to momentum-space at (c-e); (g) photoluminescence intensity (black circles), full-width at half-maximum (FWHM, red squares), and (h) energy of the lasing state versus pump power at zero voltage, extracted from the fitting of the emission spectral profile obtained by integrating dispersion data around $k_{x}=0$ over ±$\pm 0.2\mu m^{-1}$ of the energy-resolved momentum space measurements. Error bars of y-axis are extracted from the covariance matrix of the fitting, while laser shot-to-shot instability is responsible for the error in x-axis. For better visibility energy resolved images (c-e) are illustrated in logarithmic scale saturated below 0.05.
@@
Figure 1(g,h) presents the dependence of PL intensity, FWHM, and emission energy on the pumping power at 0 mV external voltage. Initially, a linear increase in intensity is observed, followed by a sharp nonlinear behaviour and a step-like drop in FWHM above the threshold pump pulse energy $P_{th}\approx 0.6$ $\mu J$, both hallmarks of a transition from spontaneous emission to photonic lasing. This transition is further evidenced by a sharp blueshift of the emission relative to the cavity dispersion minimum below threshold. In conventional inorganic systems, such as GaAs-quantum well microcavities with Wannier–Mott excitons, such blueshift near threshold is a recognized signature of polariton lasing [44]. However, this interpretation does not straightforwardly apply to organic microcavities with Frenkel excitons, which are highly localized and do not support comparable repulsive interactions. It has been shown [45, 46] that, in organic systems, the blueshift primarily arises from the saturation of molecular transitions and the associated refractive index change and hence, is present in both weak and strong light-matter coupling regimes and, importantly, is wavelength dependent, see Fig 1(c,d,e). Under tight optical pumping, it creates a localized potential sufficient to induce the outflow of coherent photons with finite average absolute value of in-plane momentum ($k_{||}\neq 0$), resulting in ring-shaped emission in momentum-space (or two split rings relative to $k_{x}=0$ in the RD regime), with the radius $k_{l}$ determined by the amount of pump-induced blueshift [see momentum-space PL at the bottom of Figs. 1(c,d,e)]. This ring is similar to resonant Rayleigh scattering ring when waves are scattering of disorder and couple to many different momenta [47]. Moreover, for tuneable LCMCs, this blueshift depends on the spectral detuning between the P580 dye absorption and the cavity mode, which can be modulated via applied voltage [Fig. 1(c,d)]. Consequently, the $k$-space profile of the coherent emission is electrically controllable through the modulation of the local potential. While the underlying mechanisms differ between room-temperature organic microcavities and cryogenically operated GaAs-based systems, both generate coherent emission with finite in-plane momentum, which is a key requirement for mutual phase-locking between lasing sites and the emergence of spatially extended lasing "supermodes" via in-plane coherent energy exchange.
@@ 2.2 Dyad lasing supermode
Figure 2: Dyad supermode lasing state. (a-i) Experimental images of the dyad supermode lasing state emission in (a-c) real-space, (d-f) momentum space, and energy-resolved momentum-space along $k_{y}=0$ (g,h,i) for two pump spots (black circles) separated by (a,d,g) 9 $\mu$m, (b,e,h) 11 $\mu$m, and (c,f,i) 13 $\mu$m. (j,k) Numerically simulated (j) real-space and (k) momentum-space distribution of the supermode lasing state emission corresponding to the experimental data at (b,e). (l) Dependence of the extended lasing state mode number n on the pump separation distance. The error bars are defined by the fragmentation of the lasing state within the pumping spot area. Each spot was pumped with the fixed spot size Gaussian beam at $P_{1,2}=1.5P^{(1)}_{th}$, where $P^{(1)}_{th}$ is the lasing threshold for a single isolated lasing state. For better visibility real-space images (a-c,j) are illustrated in logarithmic scale saturated below 0.01.
@@
Fig. 2(a-f) shows the experimentally measured real- (a-c) and momentum- (d-f) space distributions of the PL from the spatially extended supermode lasing state formed due to the mutual exchange of coherent photons out-flowing from each of the two individually pumped lasing spots. In the absence of coherent energy exchange one would observe two independent lasing spots in the real space with the PL intensity monotonically decaying outside of the pumping spot areas, Fig. 1(b), while Fourier-space would have a ring-shaped pattern similar to the single state case. Instead, here we observe phase-locking between two individually pumped lasing states, evidenced by a characteristic oscillatory interference pattern of PL intensity in both real and Fourier spaces along the axis connecting the pump spots. This interference pattern directly reveals a macroscopic coherent wavefunction extending over distances up to $d=20\mu\mathrm{m}$. The resulting dyad lasing supermodes exhibit either even or odd parity with respect to their reflection axis at $x=0$. The parity is determined by the relative phase between the lasing centres ($\Delta\phi=0$ or $\pi$), depending on the separation $d$ between the pump spots, as expected and observed previously for polariton dyads [48]. Maximum intensity at the midpoint indicates in-phase synchronisation ($\Delta\phi=0$), while a minimum of the intensity in the centre reflects anti-phase synchronisation ($\Delta\phi=\pi$). In both cases, a single mode of corresponding parity was present in energy-resolved Fourier-space PL [Fig. 2(g-i)]. We quantitatively characterise the supermode by introducing the mode number $n$ corresponding to the number of nodes between the pump spots analogous to the mode numbers used to describe the higher-order Hermite-Gaussian transverse modes. A linear increase in the mode number $n$ of the spatially extended supermode is visible when the distance between pumping spots $d$ is increased by approximately half-wavelength of the outflowing photons, that is $k_{l}d=n\pi$ where $k_{l}$ is the in-plane photon outflow wavenumber, Fig. 2(l).
@@
To theoretically reproduce the real and Fourier-space PL from the dyad lasing state, we perform numerical simulations based on a two-dimensional (2D) non-Hermitian Schrödinger equation under drive and dissipation (a simplified version of Maxwell-Bloch equations [49, 50]). Details about the theoretical model employed are presented in Methods (Sec. 4.3). In Fig. 2(j,k) we show the simulation results for the dyad lasing state separated by $d=11\>\mu\mathrm{m}$ corresponding to the experimental data presented in Fig. 2(b,e). Corresponding simulations of the lasing states for $d=9\>\mu$m and $13\>\mu$m can be found in the SI Sec. 3. The slight mismatch between the experimental and numerically obtained data is due to the local disorder morphology of the sample and is mostly present in real-space PL profiles.
@@
Amazingly, in Fig. 4(c,g) we observe revival of the dyad coupling when entering the RD SOC regime at higher voltages $V=1840$ mV. Here, mixing between two distinct cavity modes results in a dispersion relation made of two off-centred cross-circularly polarized parabolas or valleys, displaced along $k_{x}$ axis parallel to the axis of the dyad [see Eq. (1) and Fig.1(e)]. Since the bottom of each dispersion branch remains similar in energy to that of the vertically polarised cavity mode at $V=0$ mV, the wavelength-dependent PL blueshift above the threshold is unchanged compared to the initial case [Fig.1(c,e)], as is the corresponding ring radius in each valley in Fourier space [Fig.4(e)]. Here, we observe emission from two separate ballistic rings in Fourier space displaced symmetrically along $k_{x}$ with opposite circular polarisation corresponding to iso-energy curves of each paraboloid (see the inset in Fig. 4(g) for $S_{3}(\mathbf{k}_{\parallel})$ Stokes parameter distribution in momentum space). We note that due to RD SOC this regime demonstrates spin-dependent anisotropy of the supermode emission with $\langle k_{x}^{\sigma+}\rangle>0>\langle k_{x}^{\sigma-}\rangle$ and $\langle k_{y}^{\sigma+}\rangle=\langle k_{y}^{\sigma-}\rangle=0$ with a clear fringe pattern observed in Fourier space, proving the phase-locking revival between the two pumped lasing states. Phase-locking revival was also observed in restoring of spatial coherence in the interferometry measurements, see Sec 1.2 in SI. In real space the total intensity of the fringes between the pump spots is lowered relative to the case of $V=0$ mV. We attribute that to the spatial redistribution of the phase-locked dyad emission in presence of spin-orbit coupling when each of the circularly polarised components has oscillating intensity pattern spatially shifted relative to each other. It results in smoothed modulation of the total intensity, while fringe pattern is clearly observed for each circular polarisation separately, see in SI Sec 1.3.
@@
where $m_{x,y}$ are renormalized effective photon masses along the $x$ and $y$ directions [37] and $\delta$ describes splitting between horizontal and vertical linear polarized modes at $k_{\parallel}=0$. The modified dispersion leads to a higher photon population in the mode with linear polarisation component parallel to the pump polarization. The emission still partially occupies two ballistic circles in the momentum space PL as seen at high wavenumbers in Fig. 4(h). The radius of the circles is larger since the emission energy is now located further above the minimum of the dispersion (2). Interestingly, here the dyad mostly populates the low momentum intersect between the two circles in momentum space corresponding to photons with mostly linear vertical polarization projection. The result is a peculiar distribution of the emission in Fourier space in which the emission anisotropy is reversed relative to the case of pure RD SOC regime ($\langle|k_{y}|\rangle>\langle|k_{x}|\rangle$). Control of the in-plane momentum anisotropy is another important feature of our system as it potentially offers a way for the realisation of a direction-dependent coupling in two-dimensional arrays of phase-locked coherent emitters. Similar to the previous case the fringe pattern in real space is distorted and observable mostly around pumping spots. However, it is clearly visible in the Fourier space over the whole area of emission distribution. We verify our interpretation through 2D non-Hermitian Schrödinger (paraxial Maxwell-Bloch) simulations shown in Fig. 5, see Sec. 4.3 for details.
@@ 2.4 Unconventional coupling in 1D lasing supermode
In coherently coupled photonic systems the nearest-neighbour (NN) coupling dominates over the next-nearest-neighbour (NNN), including systems based on coupled lasers [51], VCSELs arrays, polaritonic and photonic lattices [3]. This is easily understandable, as the coupling strength is usually defined by the mutual overlap of the coupled states wavefunctions, which decay exponentially with the distance from each state centre. Overcoming this behaviour is of prior importance for universal applications of coherently coupled photonic arrays as all-optical simulators  [52]. The possibility of NNN coupling with polaritonic condensates in GaAs-based microcavities at cryogenic temperatures has been shown previously using distance-dependant spin-screening of NN interaction leveraging the optical spin Hall effect [53]. Same approach based on the TE-TM field is not straightforwardly applicable in organic microcavities, but it can be adapted, because the polarisation basis in these systems is defined by horizontal and vertical components, following the orientation basis of the molecular dipole moment. As we show above, the polarisation of the lasing state is then inherited from the pump polarisation and is linear as long as SOC remains small, see Sec. 2.1. As a result, the efficient coupling between the lasing spots is only possible while they share the same linear polarisation inherited from the pumps. In Fig. 6 we demonstrate the realisation of NNN coupling regime in a chain of 3 coupled lasing states, by measuring the real space PL (a-c), momentum space PL (d-f), and (g-i) real space $S_{1}$ Stokes parameter of the non-normalized Stokes vector of the extended lasing state emission [see Eq. (3)]. The corresponding numerical simulations obtained via the 2D non-Hermitian Schrödinger model (Sec. 4.3) are shown in the SI, Sec. 3.
@@
However, when we flip the polarisation of the central pump spot to the vertical one, both real and Fourier-space distributions change dramatically, see Fig. 6(c,f,i). First, the fringes in the real space are not visible, due to the low intensity similar to the case of the far-separated dyad [Fig. 6(a)]. The Fourier space shows a superposition of two states, see Sec. 1.4 of SI for details. The first originates from a NNN type of coupling between the horisontally polarised edge states which is visible with the odd parity and the same periodicity as the one observed for the far-separated dyad. The second mode is ring-shaped and corresponds to the Fourier space profile of an isolated vertically polarised central lasing state. Thus we eliminate the coupling between the nearest neighbours by exciting the lasing state in the middle with the opposite polarisation to the outer pump spots, while the NNN coupling is still present.